# Evolutionary Profiles for Protein Fitness Prediction

### v2

@@ Abstract
Predicting the fitness impact of mutations is central to protein engineering but constrained by limited assays relative to the size of sequence space. Protein language models (pLMs) trained with masked language modeling (MLM) exhibit strong zero-shot fitness prediction; we provide a unifying view by interpreting natural evolution as implicit reward maximization and MLM as inverse reinforcement learning (IRL), in which extant sequences act as expert demonstrations and pLM log-odds serve as fitness estimates. Building on this perspective, we introduce EvoIF, a lightweight model that integrates two complementary sources of evolutionary signal: (i) within-family profiles from retrieved homologs and (ii) cross-family structural–evolutionary constraints distilled from inverse folding logits. EvoIF fuses sequence–structure representations with these profiles via a compact transition block, yielding calibrated probabilities for log-odds scoring. On ProteinGym (217 mutational assays; >2.5M mutants), EvoIF and its MSA-enabled variant achieve state-of-the-art or competitive performance while using only 0.15% of the training data and fewer parameters than recent large models. Ablations confirm that within-family and cross-family profiles are complementary, improving robustness across function types, MSA depths, taxa, and mutation depths. The codes will be made publicly available at https://github.com/aim-uofa/EvoIF.
@@ 1 Introduction
Protein fitness prediction is constrained by the scarcity of experimental measurements relative to the vastness of protein space [Biswas et al., 2021]. Consequently, self-supervised methods for protein representation learning have become essential for protein fitness prediction [Hopf et al., 2017; Meier et al., 2021a; Zhang et al., 2024]. Recently, protein language models (pLMs) including ESM series [Rives et al., 2019; Lin et al., 2022] and their structure-informed variants [Hsu et al., 2022], trained through masked language modeling (MLM), have demonstrated remarkable zero-shot capabilities in protein fitness prediction [Notin et al., 2023]. These models can predict the impact of mutations on protein function without additional training specific to particular protein families, sometimes achieving performance comparable to specially trained models. Current state-of-the-art approaches, including AIDO-Protein-RAG [Li et al., 2024a] and VenusREM [Tan et al., 2024], further boost performance by integrating homologous sequences as evolutionary context.
@@
Issue 3. Existing models have not fully considered the comprehensive modeling of protein evolutionary information. For sequence evolution information, researchers have applied Multiple Sequence Alignment (MSA) [Edgar and Batzoglou, 2006] for modeling. In contrast, Inverse Folding (IF) [Hsiao et al., 2023] has been developed to model cross-family structural evolutionary information. Notably, MSA relies solely on sequences, while IF depends solely on structure. Therefore, for a protein with both sequence and structure, it is natural to construct a comprehensive evolutionary model that incorporates both its sequence and structural information. However, this aspect remains underexplored. The majority of research treats structure merely as part of protein representation, overlooking the evolutionary information embedded within it.
@@
2) We explicitly incorporate sequence evolutionary information from homologous sequences of the same family into the model. This information is obtained through sequence similarity searches [Edgar and Batzoglou, 2006], or structure similarity searches such as Foldseek [Van Kempen et al., 2024], to identify the most closely related sequences within the same family. These sequences exhibit the most direct sequence or structure homology and have been shown to be beneficial for predicting protein fitness [Li et al., 2024a; Tan et al., 2024]. This approach can be viewed as a form of in-context reinforcement learning, where homologous sequences act as supplementary expert demonstrations. By providing family-specific contextual information, these homologous sequences enhance the basis for protein fitness prediction.
@@ 2.1 Protein Language Models for Fitness Prediction
Notation and assumption. We focus on substitutions and, consistent with common practice, assume that a small number of substitutions do not alter the protein’s backbone structure [Tan et al., 2025; Su et al., 2023; Zhang et al., 2024; Li et al., 2024b; Tan et al., 2024; Sun et al., 2024; Li et al., 2024a]. Given a wild-type protein with sequence $S^{\text{wt}}$ and structure $X^{\text{wt}}$ , its mutant has a sequence $S^{\text{mt}}$ that differs from $S^{\text{wt}}$ at the mutation sites, while its backbone structure remains unchanged ($X^{\text{wt}}=X^{\text{mt}}$). The objective is to develop an unsupervised model that predicts the fitness score for each mutant, quantifying its functional change relative to the wild-type.
@@ 2.2 Protein Evolution as a Markov Decision Process
A common practice in protein fitness prediction is to supplement pLMs with evolutionary information from homologous sequences, which has been shown to further boost performance [Li et al., 2024a; Tan et al., 2024]. Similarly, in large language models, a technique called self-evolution has emerged, where models use prior problem-solving trajectories as context to improve their reasoning and agentic abilities [Fu et al., 2025; Han et al., 2025; Wen et al., 2025; Zhao et al., 2025]. This parallel suggests an intuitive explanation: just as humans learn from examples and adapt their reasoning based on relevant context, both protein language models and general language models can benefit from incorporating evolutionary trajectories as contextual demonstrations. In the protein domain, homologous sequences retrieved via sequence similarity searches [Edgar and Batzoglou, 2006] or structure-based searches [Van Kempen et al., 2024] provide evolutionary trajectories that act as expert demonstrations, constraining the solution space to biologically plausible mutations.
@@ 3.1 Experimental Settings
  • Sequence–structure hybrid models: MIF-ST [Yang et al., 2023], ProtSSN [Tan et al., 2025], SaProt [Su et al., 2023], S2F [Zhang et al., 2024], S3F [Zhang et al., 2024], ProtSST (KK=2048) [Li et al., 2024b].
  • Structure- and MSA-hybrid models: S2F-MSA [Zhang et al., 2024], S3F-MSA [Zhang et al., 2024], VenusREM [Tan et al., 2024], AIDO-Protein-RAG 16B [Sun et al., 2024; Li et al., 2024a].
@@ B.1 Protein Fitness Prediction
Structure-informed approaches leverage 3D constraints to improve robustness and biological plausibility. ProteinMPNN [Dauparas et al., 2022], MIF [Yang et al., 2023], and ESM-IF [Hsu et al., 2022] demonstrate that incorporating geometric inductive biases benefits fitness prediction, especially for structure-sensitive properties. Hybrid sequence–structure models, including ProSST [Li et al., 2024b], ProtSSN [Tan et al., 2025], and S2F/S3F [Zhang et al., 2024], further enhance accuracy in MSA-free settings. Complementarily, MSA-enhanced hybrids such as MSA Transformer [Rao et al., 2021], Tranception and TranceptEVE [Notin et al., 2022a, b] combine family-agnostic pLMs with family-specific alignment signals. Recent systems like VenusREM [Tan et al., 2024] and AIDO-Protein-RAG [Sun et al., 2024, Li et al., 2024a] highlight the value of jointly exploiting structural and evolutionary information.

### v3

@@ Abstract
Predicting the fitness impact of mutations is central to protein engineering but constrained by limited assays relative to the size of sequence space. Protein language models (pLMs) trained with masked language modeling (MLM) exhibit strong zero-shot fitness prediction; we provide a unifying view by interpreting natural evolution as implicit reward maximization and MLM as inverse reinforcement learning (IRL), in which extant sequences act as expert demonstrations and pLM log-odds serve as fitness estimates. Building on this perspective, we introduce EvoIF, a lightweight model that integrates two complementary sources of evolutionary signal: (i) within-family profiles from retrieved homologs and (ii) cross-family structural–evolutionary constraints distilled from inverse folding logits. EvoIF fuses sequence–structure representations with these profiles via a compact transition block, yielding calibrated probabilities for log-odds scoring. On ProteinGym (217 mutational assays; >2.5M mutants), EvoIF and its MSA-enabled variant achieve state-of-the-art or competitive performance while using only 0.15% of the training data and fewer parameters than recent large models. Ablations confirm that within-family and cross-family profiles are complementary, improving robustness across function types, MSA depths, taxa, and mutation depths. The codes will be made publicly available.
@@ 1 Introduction
Protein fitness prediction is constrained by the scarcity of experimental measurements relative to the vastness of protein space [Biswas et al., 2021]. Consequently, self-supervised methods for protein representation learning have become essential for protein fitness prediction [Hopf et al., 2017; Meier et al., 2021a; Zhang et al., 2024]. Recently, protein language models (pLMs) including ESM series [Rives et al., 2019; Lin et al., 2022] and their structure-informed variants [Hsu et al., 2022], trained through masked language modeling (MLM), have demonstrated remarkable zero-shot capabilities in protein fitness prediction [Notin et al., 2023]. These models can predict the impact of mutations on protein function without additional training specific to particular protein families, sometimes achieving performance comparable to specially trained models. Current state-of-the-art approaches, including AIDO-Protein-RAG [Li et al., 2024b] and VenusREM [Tan et al., 2024], further boost performance by integrating homologous sequences as evolutionary context.
@@
Issue 3. Existing models have not fully considered the comprehensive modeling of protein evolutionary information. For sequence evolution information, researchers have applied Multiple Sequence Alignment (MSA) [Edgar and Batzoglou, 2006] for modeling. In contrast, Inverse Folding (IF) has been developed to model cross-family structural evolutionary information. Notably, MSA relies solely on sequences, while IF depends solely on structure. Therefore, for a protein with both sequence and structure, it is natural to construct a comprehensive evolutionary model that incorporates both its sequence and structural information. However, this aspect remains underexplored. The majority of research treats structure merely as part of protein representation, overlooking the evolutionary information embedded within it.
@@
2) We explicitly incorporate sequence evolutionary information from homologous sequences of the same family into the model. This information is obtained through sequence similarity searches [Edgar and Batzoglou, 2006], or structure similarity searches such as Foldseek [Van Kempen et al., 2024], to identify the most closely related sequences within the same family. These sequences exhibit the most direct sequence or structure homology and have been shown to be beneficial for predicting protein fitness [Li et al., 2024b; Tan et al., 2024]. This approach can be viewed as a form of in-context reinforcement learning, where homologous sequences act as supplementary expert demonstrations. By providing family-specific contextual information, these homologous sequences enhance the basis for protein fitness prediction.
@@ 2.1 Protein Language Models for Fitness Prediction
Notation and assumption. We focus on substitutions and, consistent with common practice, assume that a small number of substitutions do not alter the protein’s backbone structure [Tan et al., 2025; Su et al., 2023; Zhang et al., 2024; Li et al., 2024a; Tan et al., 2024; Sun et al., 2024; Li et al., 2024b]. Given a wild-type protein with sequence $S^{\text{wt}}$ and structure $X^{\text{wt}}$ , its mutant has a sequence $S^{\text{mt}}$ that differs from $S^{\text{wt}}$ at the mutation sites, while its backbone structure remains unchanged ($X^{\text{wt}}=X^{\text{mt}}$). The objective is to develop an unsupervised model that predicts the fitness score for each mutant, quantifying its functional change relative to the wild-type.
@@ 2.2 Protein Evolution as a Markov Decision Process
A common practice in protein fitness prediction is to supplement pLMs with evolutionary information from homologous sequences, which has been shown to further boost performance [Li et al., 2024b; Tan et al., 2024]. Similarly, in large language models, a technique called self-evolution has emerged, where models use prior problem-solving trajectories as context to improve their reasoning and agentic abilities [Fu et al., 2025; Han et al., 2025; Wen et al., 2025; Zhao et al., 2025]. This parallel suggests an intuitive explanation: just as humans learn from examples and adapt their reasoning based on relevant context, both protein language models and general language models can benefit from incorporating evolutionary trajectories as contextual demonstrations. In the protein domain, homologous sequences retrieved via sequence similarity searches [Edgar and Batzoglou, 2006] or structure-based searches [Van Kempen et al., 2024] provide evolutionary trajectories that act as expert demonstrations, constraining the solution space to biologically plausible mutations.
@@ 3.1 Experimental Settings
  • Sequence–structure hybrid models: MIF-ST [Yang et al., 2023], ProtSSN [Tan et al., 2025], SaProt [Su et al., 2023], S2F [Zhang et al., 2024], S3F [Zhang et al., 2024], ProtSST (KK=2048) [Li et al., 2024a].
  • Structure- and MSA-hybrid models: S2F-MSA [Zhang et al., 2024], S3F-MSA [Zhang et al., 2024], VenusREM [Tan et al., 2024], AIDO-Protein-RAG 16B [Sun et al., 2024; Li et al., 2024b].
@@ B.1 Protein Fitness Prediction
Structure-informed approaches leverage 3D constraints to improve robustness and biological plausibility. ProteinMPNN [Dauparas et al., 2022], MIF [Yang et al., 2023], and ESM-IF [Hsu et al., 2022] demonstrate that incorporating geometric inductive biases benefits fitness prediction, especially for structure-sensitive properties. Hybrid sequence–structure models, including ProSST [Li et al., 2024a], ProtSSN [Tan et al., 2025], and S2F/S3F [Zhang et al., 2024], further enhance accuracy in MSA-free settings. Complementarily, MSA-enhanced hybrids such as MSA Transformer [Rao et al., 2021], Tranception and TranceptEVE [Notin et al., 2022a, b] combine family-agnostic pLMs with family-specific alignment signals. Recent systems like VenusREM [Tan et al., 2024] and AIDO-Protein-RAG [Sun et al., 2024, Li et al., 2024b] highlight the value of jointly exploiting structural and evolutionary information.